On Wednesday afternoon, your patient returns from surgery with an IV fluid order for 1000 cc every 8 hours. On Thursday morning at 8am, you assess that 600 mL of a 1L bag has been absorbed. The physician orders the remainder of that bag to infuse over the next 6 hours. You know that the IV tubing used by your unit delivers 10 gtt/mL. What will the correct rate of flow be?
Answer #1

Ans) On Wednesday, total volume- 1000 ml

On Thursdays, 600ml has been absorbed

Remainder of the bag= 1000-600= 400ml

So to infuse 400ml over next 6 hours

Time- 6hrs = 6×60 = 360

Drop factor- 10gtts/ml

Formula:

Flow rate (ml/hr) = Volume (ml) ÷ time (mins) × drop factor

Let x be the rate of flow

x= 400/360 × 10

x= 11.11 = 11ml/min

The rate of flow is 11ml/min.